1. A culture of Achromobacter xylosodoxidans was found to increase from 2 x 103 cells to 3.8 x 103 cells in 3 hours, what is its specific growth rate?
Answer #1

Initial concentration = 2 × 10^3

Final concentration = 3.8 × 10^3

Change in the concentration of bacterial cells = 3.8 × 10^3 - 2 × 10^3 = 1.8 × 10^3 cells

Time = 3 hours = 3 × 60 = 180 minutes

Specific growth rate = change in concentration/time

= 1.8 × 10^3 / 180

= 1800/180 = 10 oer minute

Therefore specific growth rate is 10 per minute. It means 10 bacteria are arising every minutes.

Please rate.